OpenStudy (anonymous):

Please Help ME! Write the slope-intercept form of the equation of the line that passes through (4, 2) with a slope of 3. A. y = 3x - 10 B. y = 3x + 4 C. y - 2 = 3 (x - 4) D. y = 4x + 3

OpenStudy (anonymous):

y-2=3(x-4) y=3(x-4)+2=3x-12+2=3x-10 so answer A,C are correct